OpenStudy (anonymous):

how to verify sin^4x-cos^4x/sinx-cosx = sinx + cosx

terenzreignz (terenzreignz):

Magic :D \[\Large \sin^4(x) - \cos^4(x) = \left[\sin^2(x) + \cos^2(x)\right] \left[\sin^2(x) -\cos^2(x)\right]\]

OpenStudy (anonymous):

i love magic! @terenzreignz

terenzreignz (terenzreignz):

Who doesn't? :3 The asker, that's who :D